OpenStudy (anonymous):

or two similar triangle, the ratio of two corresponding sides is A1:A2=1:3. Find the ratio of their areas A1:A2

OpenStudy (anonymous):

What do we know about similar triangles?

OpenStudy (anonymous):

that they can be proportional?

OpenStudy (anonymous):

Exactly. Corresponding sides are proportional. They give us the scaling proportion. What's the equation for the area of a triangle?

OpenStudy (anonymous):

not sure how to answer that because that is the question that they are giving me

OpenStudy (anonymous):

We need the equation for an area to see exactly how they relate to one another. \[A_Triangle = \frac{1}{2} Base \times Height\]

OpenStudy (anonymous):

I just need to know ratio if it is 1 to 3

OpenStudy (anonymous):

If corresponding sides are proportional, then both the base and the height are also proportional. In our case, we know that a side of triangle 1 is 3 times the size of a side from triangle 2. Using that equation, we can relate the area of one triangle to the other.

OpenStudy (anonymous):

\[A _{1}:A _{2}=1:3 \]

OpenStudy (anonymous):

\[A_1 = \frac{1}{2} B_1 \times H_1\] \[A_2 = \frac{1}{2} Base_2 \times H_2\] And we know that: \[B_1 = 3 \times B_2\] and \[H_1 = 3 \times H_2\] Plugging those into the above equation and solving so that it looks like: \[A_1 = C \times (\frac{1}{2} B_2 \times H_2)= C \times A_2\] Shows us what we want to know.

OpenStudy (anonymous):

Where C is just some number (the proportionality constant of the areas)

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

Great! What do you get for C?

OpenStudy (anonymous):

should I plug in any numbers because there is no numbers in this problem

OpenStudy (anonymous):

You'll plug in: \[B_1 = 3B_2, H_1 = 3H_2\] Since we're told this is true in the problem (A side from triangle 1 is 3 times larger than the corresponding side from triangle 2)

OpenStudy (anonymous):

are you a teacher?

OpenStudy (anonymous):

I teach physics at a university